In any triangle, the sum of the measures of the angles is 180 degrees. In triangle ABC, angle A is 4 times as large as angle B. Angle C measures 20 degrees less than angle B. Find the measure of each angle.
Answer by MArk_HeRras(6)

A/R(representing the unknown):
4x = Angle A
x = Angle B
x-20 = Angle c
E(Equation): 4x + x + x - 20=180
S(Solution):
4x + x + 20 - x = 180
Add 4x + x - x = 4x
Then, transposed -20 to the right side showing us
180 - 20 = 160
4x = 160

Now, Divide 4 to both sides like this:
4x/4 = 160/4

The Answer is x = 40
C(Checking): 4(40) = 160 =Angle A
x = 40 = Angle B
20 - 40 = -20
4(40) + 40 + 20 - 40 = 180
160 + 60 - 40 =180
220 - 40 = 180
180 = 180